# Clinicians’ Voice: Fundamental Considerations for XAI in Healthcare

### v2

@@ 4.3 Attitude towards (X)AI in healthcare
The most common response was that AI is simply more efficient than humans and can “make our life easier” (interview 3). Again, it was highlighted that AI can take over largely the tedious tasks, hence acting not necessarily as a DST but rather taking over relatively well-defined but time-consuming tasks, such as planning and writing reports.
All interviewees had mainly positive thoughts about AI-based DSTs, most of them believing that AI can help them make the best decision. As a human being, “it’s really difficult to make complex decisions and to take into account more than four or five variables” (interview 9), which is also supported by research in cognitive sciences (e.g., Miller, 1956; Cowan, 2001; Luck and Vogel, 1997; Oberauer and Kliegl, 2006). Working memory capacity has also been linked to decision-making, and several works have shown that humans tend to rely on heuristics to make decisions if cognitive demands are too high (e.g., Bröder and Schiffer, 2003; Shah and Oppenheimer, 2008). However, it is not only about how many variables clinicians need to take into account when assessing the patient’s situation, but also relating these to their domain knowledge and previous cases. For example, one interviewee pointed out that AI “has a memory of a lot more cases than I can experience and remember” (interview 2), so it can base its prediction on a much wider range of previous patients. Another was convinced that “you can treat a patient better with the data you have” (interview 5), which will “make it better” (interview 5) for both the patient and the doctor.
While interviewees seemed to agree that AI can support them in the process of making the best decision for their patients, two explicitly mentioned that AI could do so by taking a more objective stance. That is, sometimes our judgments can be clouded by emotions and especially in fields where there is frequent clinician-patient contact one may have too much of a “tunnel vision on [a] patient” (interview 3), in which cases it would be helpful to have “an objective somebody next to you” (interview 3).
@@
However, interviewees also voiced concerns that the distance between them and their patients will become larger when using AI-based DSTs. The same case may have different outcomes because of the patients’ wishes, beliefs, and preferences – “soft features” (interview 2) that are not captured in the data. Personalized treatment is only possible if you interact and connect with your patient, know their story, and incorporate their wishes. Using AI-based devices for some tasks bears the fear that the distance between doctors and their patients becomes so large that they will be dependent on these devices. This outcome would take away a large part of what drives many clinicians, which is simply the interaction with their patients. However, it is important to note that other interviewees, when asked about their opinion on this aspect, said that they are not concerned about it because regardless of where any medical information comes from, it is still them (the clinicians) who are in charge of the situation and in control of the decision. As the name suggests, a DST is merely there to give support – the decision still lies with the clinician who takes them in consultation with the patient.
@@
Bias and AI literacy are other aspects that concern clinicians. In the context of AI, bias relates to the data the model was trained on, and has been an important topic in recent research, especially after criticisms were voiced against some algorithms implemented in practice; see, for example, BBC (2018). Bias in AI exists when the model favors certain people, that is an AI-based DST may provide more accurate guidance for a certain group of patients. For example, if the dataset used for training included a higher proportion of elderly people, the model might give misleading or incorrect results for younger patients. If such systems were to be used, clinicians should be aware of these drawbacks to accurately interpret the output given by the device. Several interviewees pointed this out and referred to it as validity. In fact, five interviewees mentioned that the validation of AI-based devices is extremely important and difficult, making it hard to develop models that can be applied to all patients. Overall, interviewees in this study showed that they are aware of the issue, and further pointed out that this is especially concerning if clinicians rely too much on the AI model instead of using it complementary to their expert knowledge. This concern was voiced especially for younger generations of clinicians who are and will be developing their expertise while these systems are around, potentially making them less critical towards them.
@@
Generally, participants seem to agree that AI is the future and can solve many problems. However, is this really true? To quote one of the interviewees, AI is “kind of the Holy Grail to a lot of people” (interview 6), without considering the context. In a significant number of cases, expert knowledge and/or simple models may be just as good as complex AI models, with the added benefit of providing concrete insights into the decision-making process (Semenova et al., 2022), which is also something several interviewees acknowledged.
@@ 4.4 What’s best – feature importances, counterfactual explanations, or rule-based methods?
Comparing only feature importance scores, counterfactual explanations and rule-based models, there is a clear tendency for people to prefer feature importance scores. Interviewees generally agreed that they are intuitive, quick and easy to understand, as you can grasp them all with one glance. It is so easy that “you don’t need to be a professional to understand what’s going on here” (interview 3). While this is clearly a pro of this method, it also raises concerns – “the […] question is, is it really then that simple?” (interview 5). Especially in complex decisions that require consideration of a multitude of factors, it is almost “not believable” (interview 1). Interviewees also pointed out that they are familiar with this method of explaining the output of AI models, and hence may be biased in preferring it over the other two methods.
@@
Counterfactual explanations require a bit more time to be understood, however they can provide one with more in-depth insights. As one interviewee put it, this type of explanation is more “fluid” (interview 5), rather than static as the feature importance scores, and it shows the effect of something changing. This property could also be used to improve prognosis, which however raises the question of causality – a change in features leads to a different prediction by the model, which does not always translate to reality. Lastly, one interviewer explained that looking at counterfactuals means one is essentially looking at other patients, and not focusing on the current patient anymore. On top of that, there are endless other patients out there, so why would one look at this specific counterfactual? This statement can be linked to discussions of distance measures in technical papers on counterfactual explanations. That is, the choice of distance function influences which counterfactual is found (Guidotti, 2024). Besides, most approaches use simple distance functions (e.g., Maragno et al., 2022, 2024; Wachter et al., 2017; Kanamori et al., 2020), however ‘closeness’ in real situations may differ from mathematical definitions.
@@
Rule-based methods are similar to how decisions are made –e.g., “Of course, if you are a two year old, you will never have a mama carcinoma”, or breast cancer in layman’s terms (interview 3)–, and they’re helpful if you really want to understand the prediction. Knowing the decision mechanism makes the AI’s prediction more “believable” (interview 1), gives a sense of control, and something concrete to base your decision on, which is especially helpful in high stakes decisions. One interviewee held the opinion that such simple model should always be preferred and we should only move towards complex AI models (and with that other types of explanations) if the simple models are not enough. However, two interviewees also pointed out that if you have many rules, this may be too complicated to use on the spot. For research purposes, when you want to uncover the mechanism behind patient characteristics and medical conditions or outcomes, rule-based methods are helpful – but in clinical practice?
@@
Regardless of the type of explanation, visual explanations are preferred for being easily understood on the spot, and because it is a way to remove any potential language barriers. Another aspect interviewees raised was uncertainty of the model prediction – being shown how certain the model is about a prediction makes it more trustworthy and helps clinicians navigate the usage of the information given. This ties in with knowing more about the background of the model, such as which data were used for training and testing, how it was validated, how big the sample was, and so on, which came up in about half of the interviews.
@@ 4.5 The role of trust in AI
One third of the interviewees mentioned trust comes automatically when a device is tested, certified, and approved. This is comparable to trusting lab results and equipment without exactly knowing how they work, or when they have last been calibrated. Somewhat in line with that, trust develops over time as you use a device and experience its benefits. Some people are hesitant towards new products, known in economics and sociology as late majority and laggards (Rogers, 2003). Especially more senior clinicians seem to be reluctant to use new AI-based medical devices. One interviewee raised the point that trust is distinct from responsibility; that is, you can trust a device but in the end you have to take responsibility for your decisions, and “how are you going to feel okay with that responsibility when it’s not your own thought?” (interview 3). This adds a new angle to the discussion on trust of AI models that has not yet been sufficiently addressed in research.
@@ 5 Discussion
Based on the results obtained, we highlight the opportunities for (X)AI research, the requirements for successful implementation of (X)AI in practice, and the main warnings provided regarding (X)AI usage and development.
@@ 5.2 Requirements for successful implementation of (X)AI in practice
Most interviewees in this study have some knowledge about (X)AI, however, mainly for their own interest in the topic. As it stands momentarily, medical education in the Netherlands does not include obligatory classes on AI or AI-based medical devices, which can have significant consequences. First, understanding the background of the model, such as the training data used and how it was validated, is important to judge its performance and get a sense of trust. Second, it is often mistaken that explanations imply causality. Using AI-based tools without sufficient background knowledge makes the users prone to misinterpret the information provided. Third, some explanations may be too simplistic and may give a false sense of understanding, like feature importance plots. Instead, explanations should stimulate critical thinking so that the user can evaluate the given output themselves. Providing too simplistic explanations may hide crucial pieces of information that otherwise would raise doubts. Especially when the provided information is in line with the user’s expectations, simple explanations may increase the risk of confirmation bias. Confirmation bias has indeed been shown to exist in human-technology interaction (e.g., Wan et al., 2022; Bauer et al., 2023) and has been acknowledged in the XAI community as well (e.g., Ghassemi et al., 2021; Rudin, 2019; Cina et al., 2023). On the other hand, experienced users may disregard a simple explanation altogether as they may find it unbelievable because the true reason for a given output is likely not as simple as a feature importance plot might suggest. After all, why are we using AI if it really were this easy?
@@
However, while we should make sure to incorporate research from other fields into the development of these products to aid their implementation, people also need to experience its practical benefits over time and need to get used to using them. By now, no-one questions the lab equipment, whereas there likely was some skepticism when they were first introduced.
@@ 5.3 Warnings regarding (X)AI usage and development
AI is supposed to be the future, but is it a good idea to use it everywhere? Lately, the field receives more and more criticism for being blindly applied to any project, just because it is AI and it looks good. However, there are many situations where a much simpler model can achieve the same level of performance as a complex black-box model (Semenova et al., 2022). Especially when it comes to high stakes situations, one should first use an interpretable glass-box model which allows clinicians to understand the mechanism, and only resort to more complex AI models if they sufficiently outperform glass-box models.
@@ 6 Conclusion
One crucial takeaway is the importance of seamlessly integrating (X)AI into clinical workflows to enhance decision-making without causing disruptions. Clinicians emphasized the need for AI literacy and stressed the significance of collaboration among AI developers, healthcare professionals, and clinicians for successful adoption.
The study also delved into clinician preferences for explanation methods, with feature importance scores being favored for their simplicity and ease of comprehension. However, counterfactual explanations and rule-based methods were valued for providing deeper insights in intricate cases. Visual explanations were deemed essential by clinicians as they enhance the intuitiveness and accessibility of AI outputs, particularly in time-sensitive situations. The choice of explanation method should align with the clinical context — feature importances may be suitable for quick decisions, while decision rules or counterfactual explanations are more appropriate for detailed analysis or audits.
In conclusion, this study advocates for transparent (X)AI systems that prioritize visual clarity, multidisciplinary collaboration, thoughtful selection of explanation methods, and integration as supportive tools rather than barriers in healthcare practice based on real-world clinician input.
@@
Appendix A Additional interview material
@@
For the toy example the algorithm returns four rules, visually displayed here:
@@
Using these rules, we can make the prediction for the highlighted patient: According to rule (1) and (2), the prediction for the highlighted patient would be class 0 (not-deceased). The summed weight for these rules is 1.0 + 0.3 = 1.3. According to rule (3), it would be predicted as class 1 (deceased), with a weight of 0.7. Rule (4) is not used for our patient, since the first condition of the rule (i.e., X1 < 3) is not met by the patient who has a value of X1 = 4. Hence, we use only the first three rules. To make the final prediction, we look at the summed weights: 1.3 and 0.7 for class 0 (not-deceased) and 1 (deceased), respectively. Hence the final prediction for the patient is 0 (not- deceased).

### v3

@@ 4.3 Attitude towards (X)AI in healthcare
The most common response (5/8) was that AI is simply more efficient than humans and can “make our life easier” (interview 3). Again, it was highlighted that AI can take over largely the tedious tasks, hence acting not necessarily as a DST but rather taking over relatively well-defined but time-consuming tasks, such as planning and writing reports.
All interviewees had mainly positive thoughts about AI-based DSTs, most of them (6/8) believing that AI can help them make the best decision. As a human being, “it’s really difficult to make complex decisions and to take into account more than four or five variables” (interview 9). However, it is not only about how many variables clinicians need to take into account when assessing the patient’s situation, but also relating these to their domain knowledge and previous cases. For example, one interviewee pointed out that AI “has a memory of a lot more cases than I can experience and remember” (interview 2), so it can base its prediction on a much wider range of previous patients. Another was convinced that “you can treat a patient better with the data you have” (interview 5), which will “make it better” (interview 5) for both the patient and the doctor.
While interviewees seemed to agree that AI can support them in the process of making the best decision for their patients, two (of eight) explicitly mentioned that AI could do so by taking a more objective stance. That is, sometimes our judgments can be clouded by emotions and especially in fields where there is frequent clinician-patient contact one may have too much of a “tunnel vision on [a] patient” (interview 3), in which cases it would be helpful to have “an objective somebody next to you” (interview 3).
Link to literature. Research in the cognitive sciences has indeed shown that humans can only consider a limited amount of information at a time (e.g., Miller, 1956; Cowan, 2001; Luck and Vogel, 1997; Oberauer and Kliegl, 2006). Working memory capacity has also been linked to decision-making, and humans tend to rely on heuristics to make decisions if cognitive demands are too high (e.g., Bröder and Schiffer, 2003; Shah and Oppenheimer, 2008).
@@
However, interviewees (2/8) also voiced concerns that the distance between them and their patients will become larger when using AI-based DSTs. The same case may have different outcomes because of the patients’ wishes, beliefs, and preferences – “soft features” (interview 2) that are not captured in the data. Personalized treatment is only possible if you interact and connect with your patient, know their story, and incorporate their wishes. Using AI-based devices for some tasks bears the fear that the distance between doctors and their patients becomes so large that they will be dependent on these devices. This outcome would take away a large part of what drives many clinicians, which is simply the interaction with their patients. However, it is important to note that four other interviewees, when asked about their opinion on this aspect, said that they are not concerned about it because regardless of where any medical information comes from, it is still them (the clinicians) who are in charge of the situation and in control of the decision. As the name suggests, a DST is merely there to give support – the decision still lies with the clinician who takes them in consultation with the patient.
@@
Bias and AI literacy are other aspects that concern clinicians. In the context of AI, bias relates to the data the model was trained on, and has been an important topic in recent research, especially after criticisms were voiced against some algorithms implemented in practice [^3] An example is an AI-based recruitment tool in testing at Amazon that discriminated on the basis of sex BBC (2018). Bias in AI exists when the model favors certain people, that is an AI-based DST may provide more accurate guidance for a certain group of patients. For example, if the dataset used for training included a higher proportion of elderly people, the model might give misleading or incorrect results for younger patients. If such systems were to be used, clinicians should be aware of these drawbacks to accurately interpret the output given by the device. Several interviewees pointed this out and referred to it as validity. In fact, five (out of nine) interviewees mentioned that the validation of AI-based devices is extremely important and difficult, making it hard to develop models that can be applied to all patients. Overall, interviewees in this study showed that they are aware of the issue, and further pointed out that this is especially concerning if clinicians rely too much on the AI model instead of using it complementary to their expert knowledge. This concern was voiced especially for younger generations of clinicians who are and will be developing their expertise while these systems are around, potentially making them less critical towards them.
@@
Generally, participants seem to agree that AI is the future and can solve many problems. However, is this really true? To quote one of the interviewees, AI is “kind of the Holy Grail to a lot of people” (interview 6), without considering the context. In line with that, this participant also emphasized that we should always start with the simple models and only move to black-box models if necessary.
Link to literature. Recently, advocates of glass-box models have highlighted that simple models like (optimal) decision trees and other rule based methods perform similarly if not better to complex black-box models, especially in cases where the data is structured like in tabular data (Rudin, 2019; Semenova et al., 2022). In fact, the well-known trade-off between performance and interpretability has also been coined as a ‘myth’ (Rudin, 2019).
@@ 4.4 What’s best – feature importances, counterfactual explanations, or rule-based methods?
Comparing only feature importance scores, counterfactual explanations and rule-based models, there is a clear tendency for people to prefer feature importance scores. Interviewees generally agreed that they are intuitive, quick and easy to understand, as you can grasp them all with one glance. It is so easy that “you don’t need to be a professional to understand what’s going on here” (interview 3). While this is clearly a pro of this method, it also raises concerns – “the […] question is, is it really then that simple?” (interview 5). Especially in complex decisions that require consideration of a multitude of factors, it is almost “not believable” (interview 1). Interviewees (2/10) also pointed out that they are familiar with this method of explaining the output of AI models, and hence may be biased in preferring it over the other two methods.
@@
Counterfactual explanations require a bit more time to be understood, however they can provide one with more in-depth insights. As one interviewee put it, this type of explanation is more “fluid” (interview 5), rather than static as the feature importance scores, and it shows the effect of something changing. This property could also be used to improve prognosis, which however raises the question of causality – a change in features leads to a different prediction by the model, which does not always translate to reality. Lastly, one interviewer explained that looking at counterfactuals means one is essentially looking at other patients, and not focusing on the current patient anymore. On top of that, there are endless other patients out there, so why would one look at this specific counterfactual?
@@
Rule-based methods are similar to how decisions are made –e.g., “Of course, if you are a two year old, you will never have a mama carcinoma”, or breast cancer in layman’s terms (interview 3)–, and they’re helpful if you really want to understand the prediction. Knowing the decision mechanism makes the AI’s prediction more “believable” (interview 1), gives a sense of control, and something concrete to base your decision on, which is especially helpful in high stakes decisions. One interviewee (out of 10) held the opinion that such simple model should always be preferred and that we should only move towards complex AI models (and with that other types of explanations) if the simple models are not enough. However, two of ten interviewees also pointed out that if you have many rules, this may be too complicated to use on the spot. For research purposes, when you want to uncover the mechanism behind patient characteristics and medical conditions or outcomes, rule-based methods are helpful – but in clinical practice?
@@
Regardless of the type of explanation, visual explanations are preferred for being easily understood on the spot, and because it is a way to remove any potential language barriers. Another aspect interviewees raised was uncertainty of the model prediction – being shown how certain the model is about a prediction makes it more trustworthy and helps clinicians navigate the usage of the information given. This ties in with knowing more about the background of the model, such as which data were used for training and testing, how it was validated, how big the sample was, and so on, which came up in about half (five) of the interviews without having posed a question targeted to this.
Link to literature. Statements by interviewees about counterfactual explanations, and specifically which counterfactual is shown, can be linked to discussions of distance measures in technical papers on counterfactual explanations. That is, the choice of distance function influences which counterfactual is found (Guidotti, 2024). Besides, most approaches use simple distance functions (e.g., Maragno et al., 2022, 2024; Wachter et al., 2017; Kanamori et al., 2020), however ‘closeness’ in real situations may differ from mathematical definitions.
@@ 4.5 The role of trust in AI
About one third (3/8) of the interviewees mentioned trust comes automatically when a device is tested, certified, and approved. This is comparable to trusting lab results and equipment without exactly knowing how they work, or when they have last been calibrated. In line with that, trust develops over time as you use a device and experience its benefits. Especially more senior clinicians seem to be reluctant to use new AI-based medical devices. One interviewee raised the point that trust is distinct from responsibility; that is, you can trust a device but in the end you have to take responsibility for your decisions, and “how are you going to feel okay with that responsibility when it’s not your own thought?” (interview 3). This adds a new angle to the discussion on trust of AI models that has not yet been sufficiently addressed in research.
Link to literature. It is not unusual that adoption of new products takes time and is not adopted at the same rate by everyone, so this aspect is not specific to AI in healthcare. In economics and sociology, the group of people that are hesitant towards new products are known as late majority and laggards (Rogers, 2003).
@@ 5 Discussion
Based on the results obtained, we highlight the opportunities for (X)AI research, the requirements for successful implementation of (X)AI in practice, and the main warnings provided regarding (X)AI usage and development; see Figure 1 for an overview.
@@ 5.2 Requirements for successful implementation of (X)AI in practice
Figure 1: Overview of key takeaways formulated on the basis of the conducted interviews.
@@
As it stands momentarily, medical education in the Netherlands does not include obligatory classes on AI or AI-based medical devices, which can have significant consequences. First, understanding the background of the model, such as the training data used and how it was validated, is important to judge its performance and get a sense of trust. Second, it is often mistaken that explanations imply causality. Using AI-based tools without sufficient background knowledge makes the users prone to misinterpret the information provided. Third, some explanations may be too simplistic and may give a false sense of understanding, like feature importance plots. Instead, explanations should stimulate critical thinking so that the user can evaluate the given output themselves. Providing too simplistic explanations may hide crucial pieces of information that otherwise would raise doubts. Especially when the provided information is in line with the user’s expectations, simple explanations may increase the risk of confirmation bias. Confirmation bias has indeed been shown to exist in human-technology interaction (e.g., Wan et al., 2022; Bauer et al., 2023) and has been acknowledged in the XAI community as well (e.g., Ghassemi et al., 2021; Rudin, 2019; Cina et al., 2023). On the other hand, experienced users may disregard a simple explanation altogether as they may find it unbelievable because the true reason for a given output is likely not as simple as a feature importance plot might suggest. After all, why are we using AI if it really were this easy?
@@
However, while we should make sure to incorporate research from other fields into the development of these products to aid their implementation, people also need to experience its practical benefits over time and need to get used to using them.
@@ 5.3 Warnings regarding (X)AI usage and development
AI is supposed to be the future, but should it be used everywhere? Lately, the field receives more and more criticism for being blindly applied to any project, just because it is AI and it looks good. However, there are many situations where a much simpler model can achieve the same level of performance as a complex black-box model (Semenova et al., 2022). Especially when it comes to high stakes situations, one should first use an interpretable glass-box model which allows clinicians to understand the mechanism, and only resort to more complex AI models if they sufficiently outperform glass-box models.
@@ 6 Conclusion
One crucial takeaway is the importance of seamlessly integrating (X)AI into clinical workflows to enhance decision-making without causing disruptions. Clinicians emphasized the need for AI literacy and stressed the significance of collaboration among AI developers, healthcare professionals, and clinicians for successful adoption. The study also delved into clinician preferences for explanation methods, with feature importance scores being favored for their simplicity and ease of comprehension. However, counterfactual explanations and rule-based methods were valued for providing deeper insights in intricate cases. Visual explanations were deemed essential by clinicians as they enhance the intuitiveness and accessibility of AI outputs, particularly in time-sensitive situations. The choice of explanation method should align with the clinical context — feature importances may be suitable for quick decisions, while decision rules or counterfactual explanations are more appropriate for detailed analysis or audits.
In conclusion, this study advocates for transparent (X)AI systems that prioritize visual clarity, multidisciplinary collaboration, thoughtful selection of explanation methods, and integration as supportive tools rather than barriers in healthcare practice based on real-world clinician input. Future research directions include investigating the role of responsibility as a construct distinct from trust. Additionally, collaborative research combining cognitive sciences with XAI to investigate the role of (X)AI in decision making is important. The field would benefit from a framework outlining the design process of (X)AI-based DSTs that incorporates multiple stakeholder perspectives. This is needed to ensure that these products can indeed be implemented seamlessly into the workflow and address the needs of clinicians.
@@
Appendix A Interview methodology details
To report our study design, methodology and results we follow the guidelines published in the QOREC checklist (Tong and Sainsbury, 2007). For conciseness, we have included the main aspects in the main body of the paper and leave the remaining details to this appendix.
The interviews were conducted by the first author (female) of this paper, who pursued their PhD in Transparent Machine Learning at the University of Amsterdam at the time the study was conducted. The interviewer obtained a university degree in Psychology with explicit training and analysis of qualitative studies like this one. While one potential interviewee did not follow up to arrange the interview after expressing interest (unknown reasons), however none withdrew their consent or dropped out after the informed consent form had been signed and the interview scheduled. All interviews were conducted in English and all except for one were held online via Zoom or Microsoft Teams depending on the interviewee’s preferences. The reason being that most participants were located elsewhere in the Netherlands and online meetings were deemed most convenient. There was no-one else besides the interviewer and the interviewee present during the interviews because all meetings were recorded and field notes were not necessary as accurate transcripts could be created later from the recordings. Participants were also not asked to check the transcripts, although they were reminded at the end of the interview that they could always reach out in case they wanted to retract something they had said or would like to add more information. The supplementary material was made available to interviewees beforehand however there was no strict instruction to study it before the interview. The interview guide with prompts and questions was not shared with the interviewees. It is believed that data saturation has been reached as aspects raised by interviewees became increasingly redundant and no new angles were discussed. Participants expressed interest in the final manuscript once published and will be send a copy of the paper.
Appendix B Additional interview material
@@
For the toy example the algorithm returns four rules, visually displayed below.
@@
Using these rules, we can make the prediction for the highlighted patient: According to rule (1) and (2), the prediction for the highlighted patient would be class 0 (not-deceased). The summed weight for these rules is 1.0 + 0.3 = 1.3. According to rule (3), it would be predicted as class 1 (deceased), with a weight of 0.7. Rule (4) is not used for our patient, since the first condition of the rule (i.e., X1 ¡ 3) is not met by the patient who has a value of X1 = 4. Hence, we use only the first three rules. To make the final prediction, we look at the summed weights: 1.3 and 0.7 for class 0 (not-deceased) and 1 (deceased), respectively. Hence the final prediction for the patient is 0 (not- deceased).